Choose a person at random and give him or her an iq test. the score y has the normal distribution with mean=100 and standard deviation=15. the probability that the person chosen has an iq score higher than 120 is what?
Answer by stanbon(75887)

Choose a person at random and give him or her an iq test. the score y has the normal distribution with mean=100 and standard deviation=15. the probability that the person chosen has an iq score higher than 120 is what?
--------------
z(120) = (120-100)/15 = 1.3333..
---
P(x > 120) = P(z > 1.3333..) = 0.0912
==========================================
Cheers,
Stan H.